OpenStudy (anonymous):

f(x) = x2 - 1, find f(3). is it 8

OpenStudy (anonymous):

is it 8

OpenStudy (mary.rojas):

(3)2-1=.....ANSWER ;)

OpenStudy (mary.rojas):

You plug that 3 into the "x" in the equation. No it is not 8. 3(2)-1=?

OpenStudy (mary.rojas):

6-1=5 5.